OpenStudy (babybritt002):

1/8 +2/t=17/8t I need step by step help

OpenStudy (malijhaa):

Let's solve your equation step-by-step. 1 8 + 2 t = 17 8 t Multiply all terms by t and cancel: 1t 8 +2=( 17 8 t)(t) 1 8 t+2= 17 8 t2(Simplify both sides of the equation) 1 8 t+2− 17 8 t2= 17 8 t2− 17 8 t2(Subtract 17/8t^2 from both sides) −17 8 t2+ 1 8 t+2=0 t= −b±√b2−4ac 2a (Use quadratic formula with a=-2.125, b=0.125, c=2) t= −(0.125)±√(0.125)2−4(−2.125)(2) 2(−2.125) t= −0.125±√17.015625 −4.25 t=−0.9411764705882353,1 Check answers. (Plug them in to make sure they work.) t=−0.941176(Works in original equation) t=1(Works in original equation)

OpenStudy (malijhaa):

Answer: t=−0.941176 or t=1

alones (alones):

@malijhaa aren't you using http://mathpapa.com/algebra-calculator.html..

OpenStudy (malijhaa):

Yess

OpenStudy (calculusxy):

\(\large \frac{1}{8} + \frac{2}{t} = \frac{17}{8t}\) Find the common denominator. What would that be? (Hint: multiply the two denominators)

OpenStudy (calculusxy):

\(\large \frac{1}{8} + \frac{2}{t} = \frac{17}{8t}\) Multiply 8 \(\times\) t to get \(8t\) Make sure that each of the fractions are equivalent to have the denominator of 8t. \[\large (\frac{8t}{1} \times \frac{1}{8}) + (\frac{8t}{1} \times \frac{2}{t} ) \rightarrow \color{red}{\frac{t}{8t} + \frac{16}{8t}}\]

alones (alones):

@calculusxy explained much bette than @malijhaa who just copied and pasted it~

OpenStudy (calculusxy):

\(\large \color{red}{\frac{t}{8t} + \frac{16}{8t} = \frac{17}{8t}}\) Now since all of the them have the common denominator of 8t, then we can simply solve for t algebraically. \(\color{green}{t + 16 = 17}\) \(\color{green}{t + 16 - 16 = 17 - 16}\) \(\large \color{green}{t = 1}\)

OpenStudy (babybritt002):

@AloneS right

OpenStudy (calculusxy):

I think that this method is much easier compared to using the quadratic formula.

alones (alones):

@babybritt002 medal him ;)